# pfl-research: simulation framework for accelerating research in Private Federated Learning

### v1

@@ Abstract
Federated learning (FL) is an emerging machine learning (ML) training paradigm where clients own their data and collaborate to train a global model, without revealing any data to the server and other participants. Researchers commonly perform experiments in a simulation environment to quickly iterate on ideas. However, existing open-source tools do not offer the efficiency required to simulate FL on larger and more realistic FL datasets. We introduce pfl-research, a fast, modular, and easy-to-use Python framework for simulating FL. It supports TensorFlow, PyTorch, and non-neural network models, and is tightly integrated with state-of-the-art privacy algorithms. We study the speed of open-source FL frameworks and show that pfl-research is 7-72$\times$ faster than alternative open-source frameworks on common cross-device setups. Such speedup will significantly boost the productivity of the FL research community and enable testing hypotheses on realistic FL datasets that were previously too resource intensive. We release a suite of benchmarks that evaluates an algorithm’s overall performance on a diverse set of realistic scenarios. The code is available on GitHub at https://github.com/apple/pfl-research.
@@ 1 Introduction
Federated learning (FL) is an approach to collaboratively train a machine learning (ML) model between clients, with coordination by a central server konevcny2016federated . Participating clients can be either user devices connected to the internet (cross-device FL) or data centers of distinct institutions, which each store data of multiple users (cross-silo FL). The common constraint is that it is infeasible to upload the data to a central server for training because of privacy, bandwidth, or other compliance concerns gdpr2016 ; hipaa1996 . However, even though only statistics derived from the original data are uploaded to a central server, FL algorithms alone are not formally or practically privacy preserving melis2019exploiting ; boenisch2023curious ; geiping2020inverting . Therefore, it is essential to combine FL with techniques to preserve the privacy of users. Deployed systems often use secure aggregation and give differential privacy (DP) guarantees to achieve this AppleMLScenesDP2023 ; talwar2023samplable ; bonawitz2019towards ; xu2023federated . We call such systems private federated learning (PFL).
@@
(P)FL is a rapidly growing field googleScholarFederatedLearning2023 , and it is unrealistic for most research to be evaluated with a real-world (P)FL deployment as most researchers do not have access to a large-scale deployment. Even for researchers with such access, user experience constraints paulik2021federated would typically significantly limit such evaluation. Additionally, training with real edge devices is slow and typically cannot support the extensive hyperparameter tuning that may be needed.
@@
Therefore, testing hypotheses in simulation is essential for continued growth in (P)FL research. FL simulation is typically much more resource intensive than training a regular (non-federated) ML model due to the extra steps involved (see Algorithm 1), so it is of particular importance that FL simulators be efficient and scale well.
@@
The community has identified a comprehensive list of open problems in FL kairouz2021advances , and effective evaluation of proposed solutions requires high quality large-scale datasets representative of these challenges. Several realistic benchmarks have been proposed to facilitate reliable evaluation of FL algorithms (see Section 2). Unfortunately, the largest and most realistic open-source benchmark datasets for FL have not been widely adopted because of the resources required to use them. As we discuss in Section 4.1, the resource demand is exacerbated by slow simulators.
@@
We introduce pfl-research, a modular and easy-to-use framework for researchers to simulate FL and PFL training that is 7-72$\times$ faster than previous FL simulators. Our framework has well defined APIs that allow an FL researcher to implement their algorithms and bundle them into components that can be shared and combined with other algorithms. pfl-research supports both PyTorch paszke2019pytorch and TensorFlow abadi2016tensorflow , and it is straightforward to integrate with other ML frameworks. Additionally, scaling up with distributed simulations is seamless with Horovod sergeev2018horovod .
@@
Speed. pfl-research is much faster than other popular FL simulators (Section 4.1) since it simulates only the computation, not the topology, of federated learning. This can significantly increase both the productivity of (P)FL researchers and the quality of publications because it makes it feasible for researchers with limited resources to evaluate algorithms on larger datasets.
User-friendly distributed simulations. pfl-research makes it easy to transition from single process to distributed simulations with zero code changes. Scale-up spans multiple dimensions: number of processes, GPUs, and machines. Multiple processes can share a GPU to increase GPU utilization. Debugging, testing, and profiling are simpler compared to alternative FL frameworks.
Privacy integration. pfl-research is tightly integrated with state-of-the-art privacy accountants and mechanisms, enabling a convenient workflow for experimenting with PFL and combining it with orthogonal FL features and algorithms.
Non-gradient-descent training. pfl-research is also a suitable framework for researching federated learning with models that require training algorithms beyond gradient descent, such as some classical ML algorithms. We provide implementations of federated gradient boosted decision trees (GBDTs) and federated Gaussian mixture models.
Diverse and unified benchmarks. To get a complete picture of where an algorithm performs well, we provide setups for benchmarking algorithms in a variety of scenarios: datasets of diverse domains, IID / non-IID, no DP / Central DP. We aim to provide the same benchmarks for both PyTorch and TensorFlow so that results are comparable across frameworks.
pfl-research began as an inner-source project and has been used both in research and for modelling practical use cases granqvist2020improving ; yu2024momentum ; azam2023federated ; azam2023importance ; xu2023training ; bagdasaryan2022training ; koga2023population ; rodriguez2021enforcing ; paulik2021federated .
@@ 2 Related Work
There are multiple open-source efforts for realizing FL. Some focus on providing infrastructure for a production environment with real edge devices liu2021fate ; ibmfl2020ibm ; ekmefjord2022scalable ; ziller2021pysyft , some focus on simulations for research tensorflowfederated ; beutel2020flower ; garcia2022flute , and some provide both he2020fedml ; lai2022fedscale ; roth2022nvidia ; xie2022federatedscope ; reina2021openfl .
@@
There are pros and cons to each framework for simulation, but there is one validly comparable metric that is key to research velocity: speed. If an FL framework runs too slowly, researchers with constrained resources will not be able to test their hypotheses on realistic benchmarks, which can outweigh other framework benefits. For example, shysheya2023differentially and liu2023language used FLAIR in their experiments with the original TensorFlow Federated code [^1] Available at https://github.com/apple/ml-flair/tree/main and they both mentioned that the computational resource requirements were substantial enough for them to have to reduce the scope of their experiments. Therefore, we mainly compare previous work with pfl-research by implementing the same cross-device benchmarks and measuring wall-clock time to finish one simulation run; see Section 4.1 for the results.
@@
The existing FL simulation frameworks and additional works from liu2022unifed ; hu2022oarf ; liang2021flbench ; chai2020fedeval provide benchmarks for evaluating general FL algorithms. Additionally, terrail2022flamby focuses on evaluation of cross-silo FL, and wu2022motley ; chen2022pfl ; matsuda2023benchmark on personalized FL. A subset of each suite of benchmarks are datasets that resemble realistic FL partitions. Unfortunately, publications have largely been focused on datasets such as CIFAR10 krizhevsky2009learning , MNIST lecun1998gradient , Shakespeare mcmahan2017communication , and LibSVM chang2011libsvm that are not, in our view, generated by processes that adequately represent realistic FL settings The large datasets realistic for FL cannot be used unless simulations become much faster. For example, only two koga2023population ; ortega2023asynchronous out of 40 accepted publications in the FL workshop at ICML 2023 that provided simulation results tested their hypothesis on a dataset with more than $4000$ user partitions (irrespective of whether it was a realistic FL dataset) FLICML2023 . Appendix A.1 of beutel2020flower confirmed the same trend for previous years.
@@
Few FL frameworks integrate security and privacy features. The consensus is that FL itself does not preserve the privacy of individual clients boenisch2023curious ; geiping2020inverting : the server actor can inspect gradients from individual clients before aggregation, and information about clients’ data can be extracted from the model after deployment shokri2017membership . The latter issue is commonly addressed using central differential privacy (DP) el2022differential ; naseri2020local ; xu2023federated . The former issue requires, in addition to DP, secure aggregation bonawitz2017practical ; bell2020secure ; talwar2023samplable . Most secure aggregation schemes merely compute a sum, and therefore do not need to be simulated, but DP mechanisms need to be included in simulations to accurately reflect the effects of noisy statistics. There are several popular libraries that provide implementations of DP mechanisms to apply DP noise on generic vectors yousefpour2021opacus ; tensorflowprivacy . However, most existing FL frameworks do not provide a simple way to plug these DP mechanisms into existing FL algorithms while ensuring that the parameters for DP are consistent with the experimental setup.
@@ 3 System Design
pfl-research is built entirely in Python and bundled into one package, making installation easy on any platform. The framework is designed to be modular so that researchers are able to select a model and an algorithm, specify additional processing such as DP, and then quickly try these out for any use case.
In this section, we first define the major architectural components in pfl-research and the philosophy behind them. Then we describe the design for realizing fast distributed simulations. We defer the introduction to FL and DP to Appendix A. For further reading, Appendix B.1 provides in-depth description of how the data flow throughout the simulation and how the components interact with each other.
3.1 Components
Dataset. A FederatedDataset returns datasets of individual clients when iterated. It parameterizes how to partition, load, and pre-process the data of an individual user. There are two ways to wrap tf.data.Dataset and torch.utils.data.Dataset (from here on referred to as "native datasets"): one native dataset per user, or one native dataset as the entire FederatedDataset representation. If the datasets of clients are small enough such that several can be loaded into memory simultaneously, a native dataset can be defined on the same level as FederatedDataset, returning the full data tensors of one user when requesting the next tensors. In cross-device FL, user datasets are typically small enough to fit into memory, hence this integration is widely applicable to asynchronously loaded client datasets. If client datasets are large and loading the whole dataset at once is inefficient or requires too much memory, e.g. in large-scale cross-silo simulations, one native dataset can represent one client, with asynchronously loaded batches.
Model. A Model is an adapter for connecting a TensorFlow or PyTorch models to the simulator. Additionally, the Model class can be extended to implement non-neural-network models. pfl-research provides an implementation of federated gradient boosted decision trees (GBDTs) friedman2001greedy and federated Gaussian mixture models (GMMs).
Algorithm. The interface for implementing federated algorithms is FederatedAlgorithm. We generalize the responsibilities of an algorithm to be:
Aggregator. An Aggregator is a lightweight interface for (1) accumulating model updates from clients and (2) reducing partial aggregated model updates across worker processes. The default is accumulating with a sum, and reducing across workers with an all-reduce sum; see Section 3.2 for further explanation.
Backend. A Backend consumes the federated datasets, postprocessors (see below), and aggregator to implement the simulation control flow. Since pfl-research was released for running simulations, only the concrete backend currently included is SimulatedBackend.
Postprocessor. A Backend can invoke a series of Postprocessors. These can implement features such as compression, sparsification, and DP mechanisms by modifying local model updates as a postprocessing step, and optionally by postprocessing the aggregated model update before it is sent back to the algorithm for the central model update procedure. The server-side postprocessing steps are run in reversed order. It is essential to be mindful of the order of postprocessors. For example, DP mechanisms commonly fit best as the last postprocessor for local updates so that no further modification of the sensitivity of the model update happens after clipping, e.g. weighting.
Callback. TrainingProcessCallback provides a hook into the central training loop. A callback does not have access to the data or model update and is called only after the central model has been updated, hence it should not alter the learning in any way. Features that extend the callback interface are procedures performed on a cadence after central iterations. Some callback implementations that pfl-research provides are: fault-tolerant training procedure, evaluation on central dataset, exponential moving average of model, stopping criterion, reporting intermediate results (csv files, TensorBoard and Weights & Biases [^2] https://wandb.ai/), and profiling tools.
Hyperparameters. There are many hyperparameters to tune in (P)FL experiments, and there has been research on constructing schedules and adaptive approaches to simplify the hyperparameter tuning process andrew2021differentially ; reddi2020adaptive ; koskela2018learning . In pfl-research, many hyperparameters of local model training or the algorithm can be expressed as either simple types (constant for the experiment) or as a HyperParam (variable across iterations). At the start of each central iteration, the algorithm requests the current value of any HyperParam instances to set as the static value for that iteration only. A concrete hyperparameter class can in turn extend Postprocessor or TrainingProcessCallback to hook into the training loop and adapt its value based on arbitrary rules.
Metrics. Metrics have conflicting implementations in various FL frameworks. We define two types of metrics:
  • Per-user metrics — Each client returns the final metric, and the aggregation over clients’ results is the average of such metrics. This is more relevant when the performance of each client is of interest, e.g. in personalization and meta-learning.
We include a concrete example on the two kinds of metrics in Appendix B.2 .
@@ 3.2 Distributed simulations
Most existing FL frameworks have explicit communication steps as they aim to simulate the topology of a real-world FL system where the central server is a bottleneck. In many deployments, the central server can be adequately scaled, and it is valuable to have a faster simulation that ignores the topology. pfl-research simulates only the computation of FL. Communication is necessary only if the simulation runs with multiple workers (using Aggregator.worker_reduce). There is no dedicated coordinator or aggregator process needed with which all other workers communicate, unlike real-world FL. All worker processes are replicas, making debugging and transition from single-process to multi-worker training straightforward.
Figure 1: (a) The (very simple) architecture of distributed simulations. Each process is a replica with a different distributed context. One synchronous communication step is done to aggregate gradients and metrics. (b) Each process has a balanced queue of users to train.
@@
In each central iteration, a sampled cohort of user IDs is scheduled among the worker processes. Given an associated weight for each user, the SimulatedBackend will do a greedy optimization based on these weights to schedule users across the worker processes to minimize stragglers. Each worker process locally aggregates client model updates and metrics (while there are more clients to train for the current central iteration). Then, inter-worker aggregation is done with all-reduce. We defer more detailed discussion on the data flow to Appendix B.1.

### v2

@@ Abstract
Federated learning (FL) is an emerging machine learning (ML) training paradigm where clients own their data and collaborate to train a global model, without revealing any data to the server and other participants. Researchers commonly perform experiments in a simulation environment to quickly iterate on ideas. However, existing open-source tools do not offer the efficiency required to simulate FL on large and realistic FL datasets. We introduce pfl-research, a fast, modular, and easy-to-use Python framework for simulating FL. It supports TensorFlow, PyTorch, and non-neural network models, and is tightly integrated with state-of-the-art privacy algorithms. We study the speed of open-source FL frameworks and show that pfl-research is 7-72$\times$ faster than alternative open-source frameworks on common cross-device setups. Such speedup will significantly boost the productivity of the FL research community and enable testing hypotheses on realistic FL datasets that were previously too resource intensive. We release a suite of benchmarks that evaluates an algorithm’s overall performance on a diverse set of realistic scenarios. The code is available on GitHub at https://github.com/apple/pfl-research.
@@ 1 Introduction
Federated learning (FL) is an approach to collaboratively train a machine learning (ML) model between clients, with coordination by a central server [45]. Participating clients can be either user devices connected to the internet (cross-device FL) or data centers of distinct institutions, which each store data of multiple users (cross-silo FL). The common constraint is that it is infeasible to upload the data to a central server for training because of privacy, bandwidth, or other compliance concerns [2, 1]. However, even though only statistics derived from the original data are uploaded to a central server, FL algorithms alone are not formally or practically privacy preserving [63, 12, 29]. Therefore, it is essential to combine FL with techniques to preserve the privacy of users. Deployed systems often use secure aggregation and give differential privacy (DP) guarantees to achieve this [6, 81, 13, 89]. We call such systems private federated learning (PFL).
@@
(P)FL is a rapidly growing field [3], and it is unrealistic for most research to be evaluated with a real-world (P)FL deployment as most researchers do not have access to a large-scale deployment. Even for researchers with such access, user experience constraints [68] would typically significantly limit such evaluation. Additionally, training with real edge devices is slow and typically cannot support the extensive hyperparameter tuning that may be needed [43].
@@
Therefore, testing hypotheses in simulation is essential for continued growth in (P)FL research. FL simulation is typically much more resource intensive than training a regular (non-federated) ML model due to the extra steps involved (see Algorithm 1), thus it is of particular importance that FL simulators be efficient and scale well.
@@
The community has identified a comprehensive list of open problems in FL [41], and effective evaluation of proposed solutions requires high quality large-scale datasets representative of these challenges. Several realistic benchmarks have been proposed to facilitate reliable evaluation of FL algorithms (see Section 2). Unfortunately, the largest and most realistic open-source benchmark datasets for FL have not been widely adopted because of the resources required to use them. As we discuss in Section 4.1, the resource demand is exacerbated by slow simulators.
@@
We introduce pfl-research, a modular and easy-to-use framework for researchers to simulate FL and PFL training that is 7-72$\times$ faster than previous FL simulators. Our framework has well defined APIs that allow an FL researcher to implement their algorithms and bundle them into components that can be shared and combined with other algorithms. pfl-research supports both PyTorch [67] and TensorFlow [4], and it is straightforward to integrate with other ML frameworks. Additionally, scaling up with distributed simulations is seamless with Horovod [75], including multi-host, multi-GPU scenarios. Cluster workload management and scheduling systems such as slurm [40] can be used to enable running multiple simulation runs with different configurations in parallel, e.g. for hyperparameter tuning.
@@
Speed    pfl-research is much faster than other popular FL simulators (Section 4.1) since it simulates only the computation, not the topology, of federated learning. This can significantly increase both the productivity of (P)FL researchers and the quality of publications because it makes it feasible for researchers with limited resources to evaluate algorithms on larger datasets.
User-friendly distributed simulations    pfl-research makes it easy to transition from single process to distributed simulations with zero code changes. Scale-up spans multiple dimensions: number of processes, GPUs, and machines. Multiple processes can share a GPU to increase GPU utilization. Debugging, testing, and profiling are simpler compared to alternative FL frameworks.
Privacy integration    pfl-research is tightly integrated with state-of-the-art privacy accountants and mechanisms, enabling a convenient workflow for experimenting with PFL and combining it with orthogonal FL features and algorithms.
Non-gradient-descent training    pfl-research is also a suitable framework for researching federated learning with models that require training algorithms beyond gradient descent, such as some classical ML algorithms. We provide implementations of federated gradient boosted decision trees (GBDTs) and federated Gaussian mixture models (GMMs).
Diverse and unified benchmarks    To get a complete picture of where an algorithm performs well, we provide setups for benchmarking algorithms in a variety of scenarios: datasets of diverse domains, IID / non-IID, no DP / Central DP. We aim to provide the same benchmarks for both PyTorch and TensorFlow so that results are comparable across frameworks.
pfl-research began as an inner-source project and has been used both in research and for modelling practical use cases [33, 92, 8, 7, 88, 9, 44, 73, 68].
@@ 2 Related Work
There are multiple open-source efforts for realizing FL. Some focus on providing infrastructure for a production environment with real edge devices [57, 58, 25, 95], some focus on simulations for research [30, 11, 28], and some provide both [34, 50, 74, 87, 71].
@@
There are pros and cons to each framework for simulation, but there is one validly comparable metric that is key to research velocity: speed. If an FL framework runs too slowly, researchers with constrained resources will not be able to test their hypotheses on realistic benchmarks, which can outweigh other framework benefits. For example, [77] and [55] used FLAIR in their experiments with the original TensorFlow Federated code [^1] Available at https://github.com/apple/ml-flair/tree/main and they both mentioned that the computational resource requirements were substantial enough for them to have to reduce the scope of their experiments. Therefore, we mainly compare previous work with pfl-research by implementing the same cross-device benchmarks and measuring wall-clock time to finish one simulation run; see Section 4.1 for the results.
@@
The existing FL simulation frameworks and additional works from [56, 38, 53, 16] provide benchmarks for evaluating general FL algorithms. Additionally, [83] focuses on evaluation of cross-silo FL, and [86, 19, 59] on personalized FL. A subset of each suite of benchmarks are datasets that resemble realistic FL partitions. Unfortunately, publications have largely been focused on datasets such as CIFAR10 [48], MNIST [51], Shakespeare [60], and LibSVM [17] that are not, in our view, generated by processes that adequately represent realistic FL settings A lot of progress has been made on creating large-scale, realistic datasets for PFL, including for example [79, 15, 36, 34, 54, 50, 18]. However, the large datasets realistic for FL cannot be used unless simulations become much faster. We did a survey on most recent publications at NeurIPS 2023 [66], ICLR 2024 [39], ICML 2024 [69]. There are 162 publications that provide simulation results in federated learning. 142 out of 162 publications provide results on 1-100 clients, while only 6 out of the 162 publications consider more than 1000 clients for the empirical analysis. Appendix A.1 of [11] confirmed the same trend for previous years.
Few FL frameworks integrate security and privacy features, despite FL alone not preserving the privacy of individual clients [12, 29]. Client’s privacy can be violated by: (1) a server actor inspecting gradients from individual clients before aggregation; and (2) information about clients’ data being extracted from the model after deployment [76]. The second issue is commonly addressed using central differential privacy (DP) [26, 65, 89]. The first issue requires, in addition to DP, secure aggregation [14, 10, 81].
@@
Most secure aggregation schemes merely compute a sum, and therefore do not need to be simulated, but DP mechanisms need to be included in PFL simulations to accurately reflect the effects of noisy statistics. There are several popular libraries that provide implementations of DP mechanisms to apply DP noise on generic vectors [91, 31]. However, most existing FL frameworks do not provide a simple way to plug these DP mechanisms into existing FL algorithms whilst ensuring consistency of DP parameters with FL hyperparameters in experiments.
@@ 3 System design
pfl-research is built entirely in Python and bundled into one package, making installation easy on any platform. The framework is designed to be modular such that researchers are able to select a model and an algorithm, specify additional processing such as DP, and then quickly try these out for any use case.
In this section, we highlight key design elements that differentiate pfl-research from alternative FL frameworks. We defer the introduction to FL and DP to Appendix A, and of the detailed system design to Appendix B, with the specific classes to which these extension points map in pfl-research to Appendix B.1.
In pfl-research simulations, the algorithm drives the training procedure by: constructing queries to execute on sampled cohorts of one or multiple federated datasets for each central iteration, defining the local optimization procedure with the assistance of a local model to execute the query for a user, defining how to use aggregated statistics to update the central state, and determining when to stop training. The output of the local optimization procedure and output of the aggregation can be postprocessed by various composable features like DP mechanisms, weighting, sparsification and compression. Algorithm 1 describe this generalized PFL simulation algorithm in detail.
We highlight these key elements that contribute to the high efficiency of pfl-research:
  1. 1.
    Only one model per worker process is initialized and preserved on the GPU at all times.
  2. 2.
    Model parameters are updated in-place on the single model the worker process has access to. In practice, the same instantiation is used for both local training, local evaluation, central updates and central evaluation. It is only the model’s state that is sometimes cloned, e.g. before training first user in new sampled cohort, but it is always cloned to already allocated tensors on the GPU. There is no memory in the order of the model size that is released and re-allocated during the course of the FL simulation.
  3. 3.
    pfl-research does not simulate the topology of FL, i.e. there is no main coordinating process to gather and broadcast model updates. Further details in Section 3.1.
  4. 4.
    ML framework tensors are used on the GPU end-to-end in the outer training loop. This includes DP mechanisms. Alternative frameworks, e.g. Flower, use NumPy for parts of the outer loop.
  5. 5.
    pfl-research performs load balancing of training users. This is particularly important in the case of a high dispersion in user dataset lengths, e.g. FLAIR dataset. A detailed analysis on the effect of this feature is presented in Appendix B.6.
  6. 6.
    Using torch.utils.data and tf.data we provide support for asynchronously loading and preprocessing entire user datasets.
pfl-research specifically emphasizes the flexibility for experimenting with DP. The framework is superior to existing (P)FL frameworks for simulating FL with privacy guarantees in several ways: (1) Central and local DP mechanisms and accounting methods are abstracted as pluggable components, allowing researchers to easily experiment with different combinations of DP methods; (2) Tight integration between the DP mechanisms and FL hyperparameters to prevent errors in using DP with FL, e.g. DP noise is correctly scaled according to the actual clipping bound used in each iteration; (3) DP mechanisms are implemented with GPU acceleration without data transferring between CPU and GPU; (4) pfl-researchoffers a broad range of state-of-the-art privacy mechanisms and accountants out-of-the-box for different use-cases, listed in Appendix B.5.
pfl-research provides interfaces that allow straightforward integration with Foundation Model (FM) frameworks such as HuggingFace [85]. The open-sourced models and datasets for FM can be easily adapted and trained in pfl-research, which we believe will enable and accelerate the frontier research on FM with PFL.
@@ 3.1 Distributed simulations
Most existing FL frameworks have explicit communication steps as they aim to simulate the topology of a real-world FL system where the central server is a bottleneck. In many deployments, the central server can be adequately scaled, and it is valuable to have a faster simulation that ignores the topology. pfl-research simulates only the computation of FL. Communication is necessary only if the simulation runs with multiple workers (using Aggregator.worker_reduce, see Appendix B.2). There is no dedicated coordinator or aggregator process needed with which all other workers communicate, unlike real-world FL. All worker processes are replicas, making debugging and transition from single-process to multi-worker training straightforward.
Figure 1: (a) The (simplified) architecture of distributed simulations. Each process is a replica with a different distributed context. One synchronous communication step is done to aggregate gradients and metrics. (b) Each process has a balanced queue of users to train.
@@
In each central iteration, a sampled cohort of user IDs is scheduled among the worker processes. Given an associated weight for each user, pfl-research will do a greedy optimization based on these weights to schedule users across the worker processes to minimize stragglers. See Appendix B.6 for further details on the scheduling algorithm. Each worker process locally aggregates client model updates and metrics (while there are more clients to train for the current central iteration). Then, inter-worker aggregation is done with all-reduce.